OpenStudy (anonymous):

If the local linear approximation of f(x) = 2cos x + e2x at x = 2 is used to find the approximation for f(2.1), then the % error of this approximation is greater than 15% between 11% and 15% between 5% and 10% between 0% and 4

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

how far did you get here?

OpenStudy (anonymous):

well i found the linear approximation for f(2.1) using x = 2

jimthompson5910 (jim_thompson5910):

what did you get for f ' (x)

OpenStudy (anonymous):

tell me if i did thi right f(2.1) = 53.76 + 107.38(x - 2) f(2.1) = 107.38x - 161

jimthompson5910 (jim_thompson5910):

I'm getting y = 107.37771x - 160.98955 for the tangent line

jimthompson5910 (jim_thompson5910):

let g(x) = 107.37771x - 160.98955 compute g(2.1) and compare that with f(2.1)

OpenStudy (anonymous):

g(2.1) = 64.503641 f(2.1) = 65.67663883

jimthompson5910 (jim_thompson5910):

now compute the percent error using those f(2.1) and g(2.1) values

OpenStudy (anonymous):

so just divide them by each other?

jimthompson5910 (jim_thompson5910):

\[\Large \text{Percent Error} = \frac{|\text{Approximate Value}-\text{Exact Value}|}{\text{Exact Value}}\] \[\Large \text{Percent Error} = \frac{|g(2.1)-f(2.1)|}{f(2.1)}\] \[\Large \text{Percent Error} = ??\]

OpenStudy (anonymous):

1.81

OpenStudy (anonymous):

are you sure it's not backwards on the numerator

jimthompson5910 (jim_thompson5910):

the order in the numerator does not matter because |x - y| = |y - x|

jimthompson5910 (jim_thompson5910):

if the absolute value bars weren't there, then yeah, the order matters

OpenStudy (anonymous):

okay well thank you very much :D!!!

jimthompson5910 (jim_thompson5910):

I'm getting roughly 0.01786 which is approximately 1.786% error

OpenStudy (anonymous):

yeah me too either way its D

jimthompson5910 (jim_thompson5910):

yes